OpenStudy (anonymous):

HELP!!! Solve using elimination. 2m+3n=12 -5m+n=-13 I know how to do it but I'm just getting the weirdest answer... I got m = -27/17 and n= 86/17. Help would be GREATLY appreciated!

OpenStudy (anonymous):

2m + 3n = 12 -5m + n = -13 ---------------- Multiplying each side by 3 on the second equation... 2m + 3n = 12 -15m + 3n = -39 ---------------- Subtracting both equations... 17m = 51 Try it from here

OpenStudy (anonymous):

Did you not get it?

OpenStudy (anonymous):

@chandhooguy You'd have to multiply by -3 because 3n has to cancel out by it's opposite :)

OpenStudy (anonymous):

ddw, I subtracted each side, so I canceled them out when I did that.

OpenStudy (anonymous):

@chandhooguy My teacher taught me by adding

OpenStudy (anonymous):

Ok, so multiply the 2nd equation by -1 and see what you get.